# PARAPHRASUS: A Comprehensive Benchmark for Evaluating Paraphrase Detection Models

### v1

@@ 6.1 Can we improve the XLM-R training?
In the main results, we observed that fine-tuning a model on the pawsx training set yields strong performance on the pawsx test set. However, despite this success, the trained model ranks last overall on paraphrasus. A natural question then is: Why did the model fail to generalize?
Is it the adversarial character? We hypothesize that this may be due to the adversarial character of pawsx. While its difficulty should ideally teach the model to handle even the most challenging paraphrase cases, the artificial construction of the dataset may have introduced biases. These biases could have led the model to rely on spurious correlations—similar to the "Clever Hans" effect — where the model learns to exploit patterns that do not generalize well, potentially explaining its struggles with broader generalization (Niven and Kao, 2019).
Ablation study A: Adding easy negatives to the training data. To test this hypothesis, we introduce varying amounts of easy negative samples into the training data. To maintain the original label distribution, we generate these easy negatives by replacing one part of an adversarial negative pair with a random sentence from the training set. The results are presented in Table 3. We observe that this strategy can indeed improve the generalization ability of the trained model, but only when the amount of easy negative samples is kept in check. Adding a moderate 25% of easy negatives reduces the model’s overall error from 27.5 to 26.3, with substantial improvements on sick (-4.3 pp) and xnli (-5.9 pp).
Is it the labels? XLM-R models trained on pawsx exhibit poor generalization on paraphrasus, which could be attributed to the target labels assigned to the synthetic sentence pairs.
Ablation Study B: Relabeling the Training Data with LLama 3
To examine whether the model’s generalization ability could be improved by utilizing labels generated by large language models (LLMs), we trained the model on the PAWS-Wiki training set using silver labels predicted by LLama 3 through zero-shot prompting. Contrary to our hypothesis, the results in Table 4, show that models trained on these silver labels exhibit even poorer generalization performance, invalidating the assumption that relabeling enhances generalization.
@@
What is the underlying cause of the model’s lack of generalization? Based on our experiments, we cannot provide a definitive answer. The first experiment indicates that incorporating simple negative samples and reducing the adversarial nature of the training data may improve generalization. This adjustment also exposes the model to a broader spectrum of semantic similarity between sentence pairs, as opposed to the original dataset, which consisted primarily of difficult positives and negatives. Therefore, we conclude that good performance on pawsx does not always mean strong generalization in paraphrase detection overall.
@@ 6.2 Human Paraphrase Understanding Study
Figure 3: Cohens $\kappa$ between humans and systems when annotating the most similar sentences from STS.
@@
We calculate pairwise Cohen’s Kappa inter-annotator agreement scores Cohen (1960) and show the results in Figure 3. We see that there is moderate to high agreement within humans and each model type. Then there is low agreement of LLama with human, and basically zero agreement of XLM with both human and LLama. Interestingly, when taking the STS maximum label as paraphrase as is intended by the task, we see that both humans do only moderately agree with this annotation. Occasionally, both humans assigned a paraphrase to a pair from STS that has a score lower of 5.0, at other times, both humans agreed to assign a non-paraphrase label to an STS pair labeled with 5.0.
@@ 6.3 Failure modes of LLM
Table 5: Examples of LLM failures on NLI and AMR-repurposed data. In the first three pairs, all LLama 3 prompts (P1, P2, and P3) incorrectly classify the pairs as paraphrases, even though they are not. In contrast, in the last three pairs, the prompts fail to assign a paraphrase label, although all examples are indeed paraphrases
When analyzing LLM performance on paraphrasus using datasets containing only certified negatives or certified positives, we observed that while the error rate of the LLMs was low, it was never zero. Table 5 shows three pairs each from the repurposed NLI xnli dataset (non-paraphrases only), as well as from the AMR-guidelines repurposed data (paraphrases only). In all these cases, all LLM prompts agree – on the wrong decision. Interestingly, sometimes, the LLM seems tricked by seemingly simple active/passive variants and assigned a negative prediction, like X saddened Y / Y was saddened by X or a slight positional variation of an adverbial particle (look up X / look X up). At times, the model mistakenly labeled a contradiction as a paraphrase, as seen in the first pair where one sentence calls something a significant threat, while the other denies it was a threat.
Overall, we should note that LLMs generally show low error rates on the NLI-repurposed data, demonstrating robustness against hard negatives. However, the error rate is higher on hard positives in the AMR-repurposed data. In 92% of cases where one or more LLM prompts make an error, at least one prompt still makes the correct decision. Despite this, our analysis indicates that LLMs can still be misled by relatively simple linguistic variations.
@@ 7 Conclusion
A key contribution of our work is the introduction of the paraphrasus benchmark, which includes sentence pairs from ten domains and covers three paraphrase detection tasks. We are releasing this benchmark to the research community, with the goal of facilitating further exploration of paraphrase phenomena, evaluating paraphrase detection models and zero-shot capabilities of LLMs, and pinpointing specific areas for improvement.
Our experiments using paraphrasus revealed several interesting insights: 1. None of the tested LLM and classifier setups demonstrated consistently strong performance across the full spectrum of paraphrases captured by the benchmark, highlighting its objective nature and the need for continued system development. 2. Even advanced LLMs as LLama 3 failed to accurately and consistently detect paraphrases – in some cases, only the passivization of the verb is all it takes to confuse all tested prompt variants. 3. Improving training strategies for smaller, more efficient models is challenging, but not impossible. Specifically, carefully inserting a certain amount of easy negatives into a large training set of adversarial pairs helped to lower the average error. 4. Among all tested LLM prompting strategies, each exhibited specific strengths and weaknesses depending on the particular tasks within the paraphrasus. However, the simplest prompt – directly asking whether the sentence pairs are paraphrases – performed best.
@@
Our proposed dataset includes text pairs with different levels of semantic and lexical similarity from diverse domains and covers many interesting paraphrasing phenomena. Does it cover all possible phenomena related to paraphrases and paraphrasing? Clearly not. For example, an obvious limitation is its coverage of phenomena in languages other than English. While our dataset does include a total of 25,360 non-English samples they aren’t spread across all ten parts of paraphrasus (e.g., the AMR-sourced dataset is only English). Thus, expanding paraphrasus to include more datasets in languages other than English, particularly low-resource languages, appears as a fruitful avenue for future work. We also invite the help of the community for this, and will allow the option to collaboratively work on extending paraphrasus.
@@ Acknowledgements
This research is conducted under the project Impresso - Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
@@ Appendix A ICL Prompt
In Context Learning (k=4) Paraphrase Detection Prompt Here are examples sentence pairs that are {paraphrase notion} (Yes) and not {paraphrase notion} (No):
@@
Figure 4: In Context Learning Prompt Template. For P1, P2, and P3, the questions asked are “paraphrases,” “semantically equivalent,” and “expressing the same content,” respectively.
Appendix B Additional Datasets Statistics
Figure 5: Average sentence lengths. Left plot shows symmetric datasets and right plot shows the asymmetric NLI datasets.

### v2

@@ 6.1 Can we improve the XLM-R training?
In the main results, we observed that fine-tuning a model on the pawsx training set yields strong performance on the pawsx test set. However, despite this success, the trained model ranks second to last overall on paraphrasus. A natural question then is: Why did the model fail to generalize?
Is it the adversarial character? We hypothesize that this may be due to the adversarial ch-aracter of pawsx. While its difficulty should ideally teach the model to handle even the most challenging cases, the artificial creation of the dataset may have introduced biases. These biases could have led the model to rely on spurious correlations – similar to the "Clever Hans" effect – where the model learns to exploit patterns that do not generalize well, potentially explaining its struggles with broader generalization (Niven and Kao, 2019).
Ablation study A: Adding easy negatives to the training data. To test this hypothesis, we introduce varying amounts of easy negative samples into the training data. To maintain the original label distribution, we generate these easy negatives by replacing one part of an adversarial negative pair with a random sentence from the training set. The results are presented in Table 3. We observe that this strategy can indeed improve the generalization ability of the trained model, but only when the amount of easy negative samples are up to 50%. Adding 25% of easy negatives reduces the model’s overall error from 27.5 to 26.3, with improvements on sick (-4.3 pp) and xnli (-5.9 pp).
Is it the labels? XLM-R models trained on pawsx exhibit poor generalization on paraphrasus, which could be attributed to the target labels assigned to the adversarial sentence pairs.
Ablation Study B: Relabeling the Training Data with Llama3.
To examine whether the model’s generalization ability could be improved by utilizing labels generated by large language models (LLMs), we trained the model on the PAWS-Wiki training set using silver labels predicted by Llama3 through zero-shot prompting. Contrary to our hypothesis, the results in Table 4 show that models trained on these silver labels exhibit even poorer generalization performance, invalidating the assumption that relabeling enhances generalization. We eyeballed a handful of examples where all prompts agreed on the label, but disagreed with the original one. Some of these examples suggest annotation errors, such as the following: S1: In 1923, there were 568 wz.1902 guns in the Polish inventory S2: In 1923, 568 Polish guns were in the inventory of wz.1902. In this example, PAWS assigns a negative label, while all prompts assign a positive label. The consensus seems to suggest that the LLM has realized that ‘wz.1902’ is a type of Polish gun, and all numbers do perfectly match. Though to be perfectly fair, an extremely strict interpretation could argue, e.g., that some of the guns in $S1$ might have been (e.g., accidentally) put in another inventory, hence supporting the assigned label, an insight that seems to “post-hoc” support our initial hypothesis in this experiment. But even when changing only those labels to positive where all prompts agreed on, the average error rises by about 6.8 pp. to 34.3% (see Table 4, last row).
@@
What is the underlying cause of the model’s lack of generalization? Based on our experiments, we cannot provide a definitive answer. The first experiment indicates that incorporating simple negative samples and reducing the adversarial nature of the training data may improve generalization. This adjustment also exposes the model to a broader spectrum of semantic similarity between sentence pairs, as opposed to the original dataset, which consisted primarily of difficult positives and negatives. On the other hand, the second experiment seems to indicate that the given PAWS labels are of mostly good quality. While apparently a few annotation mistakes can be corrected by LLM relabeling, any relabeling overall worsens a trained model’s generalization.
We conclude that good performance on pawsx does not always mean strong generalization in paraphrase detection overall. We have not yet identified a method to train an efficient model on PAWS that effectively generalizes to out-of-distribution scenarios. We speculate that over-fitting to complex adversarial data creation schemes may be at play, counteracting any potentially valuable learning signal (for now).
@@ 6.2 Human Paraphrase Understanding Study
Figure 3: Cohens $\kappa$ between humans and systems when annotating the sts-h dataset that consists of highly similar (STS Score 4-5) sentences.
@@
We calculate pairwise Cohen’s Kappa inter-annotator agreement scores Cohen (1960) and show the results in Figure 3. We see that there is moderate to high agreement within humans and each model type. Then there is low agreement of Llama with human, and basically zero agreement of XLM with both human and Llama. Interestingly, when considering the maximum STS label of 5 as indicative of a paraphrase, as intended by the task, we observe that human annotators exhibit only moderate agreement with this classification. Occasionally, both humans assigned a paraphrase to a pair from STS that has a score lower of 5.0, at other times, both humans agreed to assign a non-paraphrase label to an STS pair labeled with 5.0.
@@ 6.3 Failure modes of LLM
When analyzing LLM performance on paraphrasus using datasets containing only certified negatives or certified positives, we observed that while the error rate of the LLMs was low, it was never zero. Table 5 shows three pairs each from the repurposed NLI xnli dataset (non-paraphrases only), as well as from the AMR-guidelines repurposed data (paraphrases only). In all these cases, all LLM prompts agree – on the wrong decision. Interestingly, sometimes, the LLM seems to be tricked by seemingly simple active/passive variants and assigned a negative prediction, like X saddened Y / Y was saddened by X or a slight positional variation of an adverbial particle (look up X / look X up). At times, the model mistakenly labeled a contradiction as a paraphrase, as seen in the first pair where one sentence calls something a significant threat, while the other denies it was a threat.
Overall, we should note that LLMs generally show low error rates on the NLI-repurposed data, demonstrating robustness against hard negatives. However, the error rate is higher on hard positives in the AMR-repurposed data. In 92% of the cases where one or more LLMs make an error, at least one prompt still makes the correct decision. Nevertheless, our analysis indicates that LLMs can be misled by relatively simple linguistic variations.
@@
Table 5: Examples of LLM failures on NLI and AMR-repurposed data. In the first three pairs, all Llama3 prompts (P1, P2, and P3) incorrectly classify the pairs as paraphrases, even though they are not. In contrast, in the last three pairs, the prompts fail to assign a paraphrase label, although all examples are indeed paraphrases.
@@ 7 Conclusion
A key contribution of our work is the introduction of the paraphrasus benchmark, which includes sentence pairs from ten domains and covers three paraphrase detection challenges. We are releasing this benchmark to the research community, with the goal of facilitating further exploration of paraphrase phenomena, evaluating paraphrase detection models and zero-shot capabilities of LLMs, and pinpointing specific areas for improvement.
Our experiments using paraphrasus revealed several interesting insights:
  1. 1.
    None of the tested LLM and classifier setups demonstrated strong performance across the full spectrum of paraphrases captured by the benchmark, highlighting its objective nature and the need for system development.
  2. 2.
    Even modern LLMs as Llama3 failed to accurately and consistently detect paraphrases – in some cases, the passivization of the verb is all it takes to confuse all tested prompt variants.
  3. 3.
    Improving training strategies for smaller, more efficient models is challenging, but not impossible. Specifically, carefully inserting a certain amount of easy negatives into a large training set of adversarial pairs helped to lower the average error.
  4. 4.
    Among all tested LLM prompting strategies, each exhibited strengths and weaknesses depending on the particular challenge within paraphrasus. However, the simplest prompt – directly asking whether the sentence pairs are paraphrases – performed best with an average error of 20.9% across the benchmark.
@@
Our proposed benchmark includes text pairs with different levels of semantic and lexical similarity from diverse domains and covers many interesting paraphrasing phenomena. Does it cover all possible phenomena related to paraphrases and paraphrasing? Clearly not. For example, an obvious limitation is its coverage of phenomena in languages other than English. While our dataset does include a total of 25,360 non-English samples, they are not spread across all ten parts of paraphrasus (e.g., the AMR-sourced dataset is only English). Thus, expanding paraphrasus to include more datasets in languages other than English, particularly low-resource languages, appears as a fruitful avenue for future work. We invite community collaboration to help extend paraphrasus.
@@ Acknowledgements
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
@@ Appendix A ICL Prompt
In-Context-Learning (k=4) Paraphrase Detection Prompt Here are example sentence pairs that are {paraphrase notion} (Yes) and not {paraphrase notion} (No):
@@
Figure 4: In-Context-Learning Prompt Template. For P1, P2, and P3, the questions asked are “paraphrases,” “semantically equivalent,” and “expressing the same content,” respectively.
Appendix B Per Language Performance
Model en (E%) de (E%) fr (E%) es (E%) ja (E%) ko (E%) zh (E%)
[…]
Llama3 Instruct P1 34.75 43.90 43.30 43.60 49.70 50.20 47.25
Llama3 Instruct P2 33.60 37.55 39.45 38.35 45.95 46.55 43.65
Llama3 Instruct P3 29.65 35.00 35.85 34.90 45.00 45.15 40.85
Llama3 Instruct P1-ICL_K4 35.50 37.65 38.40 39.10 40.25 43.50 38.25
Llama3 Instruct P2-ICL_K4 29.90 31.20 33.05 33.85 37.45 38.70 34.70
Llama3 Instruct P3-ICL_K4 27.95 31.15 32.20 32.85 36.55 38.50 33.30
Table 6: Error rates of Table 1 for the pawsx test set per language.
Model en (E%) de (E%) fr (E%) es (E%) zh (E%)
[…]
Llama3 Instruct P1 15.18 12.71 13.16 11.87 8.74
Llama3 Instruct P2 2.23 1.44 1.18 1.15 0.79
Llama3 Instruct P3 2.40 1.24 1.14 1.20 0.63
Llama3 Instruct P1-ICL_K4 4.82 3.10 2.60 1.98 1.44
Llama3 Instruct P2-ICL_K4 1.00 0.19 0.15 0.12 0.04
Llama3 Instruct P3-ICL_K4 0.78 0.13 0.15 0.13 0.06
Table 7: Error rates of Table 1 for the XNLI subset dataset for the five languages we use.
Appendix C Additional Datasets Statistics
@@
Refer to caption A
Figure 6: Average sentence lengths. Left plot shows symmetric datasets and right plot shows the asymmetric NLI datasets of paraphrasus